Find the equivalent resistance between a and b,

(i.e., Req = Rab). Answer in units of ?.

Find the equivalent resistance between a and b, (i
Answer #1

solution gr za a im Fig Gruit diagram, a that it is unbalanced from figure it is clear Wheatstone bridge. Hence to find equivfigure is also equivalet Thus circuit to following diagram of diagram 1 in а ом- are from figure it is clear that are & in in= 7 4 + + - t = 49+66 7866 7x66 - R = 462 - R=4.02 This is equivalent resisface of R&RL. New circuit diagram- R-4.02r. series